# EdgeAlpha: Distributed Process Discovery at

### v3

@@ V-F Number of Queried Nodes
Figure 4: Comparison of average queried nodes with MFP Requesting and querying all nodes. Diagram also shows standard error. EdgeAlpha can reduce communication by querying the most frequent predecessors. With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
@@
Figure 5: Fitness over time. Intermediate result FMs of the data sets quickly converge into the FM of the traditional Alpha Miner. The BPIC 2017 data set, for example, already has a fitness value of over 90% after 200 events compared with the Alpha Miner FM, which is approx. 0.02% of the events in the event log.
@@
Figure 6: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence. A steeper curve indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
Figure 7: Relationship between batch size and average number of queried nodes per event. Normalizing the average queried nodes for comparison. Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50. With increasing batch size, the reduction rate decreases.
@@
Figure 8: Stabilization of queried node count over time in Hospital Log. Since the data fluctuates, we use moving averages to visualize the stabilization. A reference line (Query All) indicates the maximum number of requested nodes per event. The moving average stabilizes after fewer than 3,000 events. We plot an additional line with a larger window size to show the trend of the stabilization. The plots are limited to the first 6,000 events.
@@
Figure 9: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log and highlights the start activities (S). We plot the number of requested nodes when querying all nodes and the average number of queried nodes over all event. The diagram shows that activities with the highest average queries occur the least, excluding start activities.
V-F1 Comparison Across Datasets
@@
Fig. 4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor. Query All denotes the naive way to message every node in the network. This serves as a baseline to compare to. With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event. We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant. Table I shows the significant reduction in the number of queried nodes. This reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
@@
Fig. 7 shows Cumulative Distribution Functions (CDF) for all datasets. The Hospital Log has the steepest increase with over 90% of events needing queries to no more than 5% of the nodes. Conversely, the Traffic Fine dataset has few instances where the predecessor can be identified after querying just 5% of the nodes. This deviation stems from the limited number of 11 activities in the Traffic Fine dataset. Hence, within the 5% threshold, only self-loops are considered.
V-F2 Stabilization Analysis
@@
We now delve deeper into the specifics of the Hospital Log. Fig. 8 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events. There is still a higher amplitude after 5,000 events. However, there are no long-lasting deviations. This event log is a good example of stabilization due to its large activity set. Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
V-F3 Activity-wise Analysis
@@
Fig. 9 takes a closer look at the specifics of the Sepsis Cases log. Depending on the activity, the Sepsis Cases log shows varied querying patterns. In contrast to other event logs, the Sepsis Cases do not have one designated start event but six different ones. However, about 95% of cases start with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others. Although ”Release E” shows an increased query count, it is not a start activity. The increased query count results from the limited occurrence (only six times), indicating ongoing stabilization. The situation is similar for the four activities to its left. They occur infrequently, which also explains the higher standard error.
V-F4 Batch Size Analysis
@@
Fig. 7 illustrates the reduction in the average number of queried nodes per event as the batch size increases. The rate in which the number of queries decreases varies between the datasets. However, for all datasets the rate decreases with a growing batch size. For all datasets, the average number of queried nodes per event already lies under 2.5% of nodes in the network with a batch size of 40. For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log, those are on average 15.6 nodes.
Overall, our results indicate that the reduction in the number of queries depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors per activity. The experiments show that it takes the nodes some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes. While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes. One such adaptation involves employing a smaller sliding window, where nodes track activity predecessors only within a fixed window. This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
V-G Footprint Matrix Sizes
@@
In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages. These messages include the node’s local FM along with two additional Booleans indicating whether a node recorded the start or end activity of a trace. Thus, the scalability of these messages depends highly on the size of each partial FM. In EdgeAlpha, each node only keeps track of their predecessors in its FM, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities. Even for datasets like the Hospital Log, which has an activity count of 624, each node transmits only a list consisting of at most 624 elements. Compared to existing distributed Alpha Miners [19, 9], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size. Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources. For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
V-H Evaluation Discussion
@@
EdgeAlpha leverages data locality to enhance its efficiency. Unlike existing distributed Alpha Miners [19, 9], which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source. As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm. Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events. Thus, building a process model becomes substantially faster. This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
VI Conclusion
@@
In this paper, we introduce EdgeAlpha, a novel distributed approach tailored to process discovery in IoT networks. By distributing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches. We discuss how EdgeAlpha’s design principles can be applied to other process mining algorithms that initially construct footprint matrices or record direct successions, such as the Heuristics Miner [14]. Moreover, integrating Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network. The batching of predecessor queries further minimizes communication overhead. Empirical evaluations showcase EdgeAlpha’s immense reduction of messages in real-world scenarios compared to a naive predecessor determination approach. EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event. By batching predecessor queries in groups of 40 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes. We analytically and experimentally demonstrate the correctness of the algorithm and additionally show how the intermediate results of EdgeAlpha quickly converge to the final result of the Alpha Miner. We highlight how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities while still producing the same results as the Alpha Miner. By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.

### v4

@@ 5.4. Number of Queried Nodes
[Comparison of average queried nodes with MFP Requesting and querying all nodes]Average number of nodes queried with and without MFP Requesting, including standard error. MFP Requesting and knowledge of the start events reduce communication demands by a factor of 7.5 to 30 depending on the dataset.
@@
Figure 5. Fitness over time. Intermediate FMs quickly converge to the centralized computed FM. The BPIC 2017 dataset, for example, already has a fitness of over 90% after 200 events.
[…]
[Fitness over time.]Intermediate FMs quickly converge to the centralized computed FM. The BPIC 2017 data set, for example, already has a fitness of over 90% after 200 events.
@@
Figure 6. Cumulative Distribution Functions: A steeper curve indicates a higher proportion of events that queried few nodes to identify their predecessors, reflecting greater efficiency.
[…]
[Cumulative Distribution Functions]Cumulative Distribution Functions: A steeper curve indicates a higher proportion of events that queried few nodes to identify their predecessors, reflecting greater efficiency.
Figure 7. Relationship between batch size and average number of queried nodes per event, normalized. Batching efficiently reduces the number of queries.
[…]
[Relationship between batch size and average number of queried nodes per event]Relationship between batch size and average number of queried nodes per event, normalized. Batching efficiently reduces the number of quries.
@@
Figure 8. Number of nodes queried over time in the Hospital Log (moving average; green line: window size 50, step 1; orange line: window size 200, step 100). The moving average stabilizes after around 3,000 events. A reference line (Query All) indicates the maximum number of requested nodes per event. Note that the plots are limited to the first 10,000 of over 150,000 events.
[…]
[Number of nodes queried over time in the Hospital Log]Number of nodes queried over time in the Hospital Log (moving average; green line: window size 50, step 1; orange line: window size 200, step 100). The moving average stabilizes after around 3,000 events. A reference line (Query All) indicates the maximum number of requested nodes per event. Note that the plots are limited to the first 10,000 of over 150,000 events.
@@
Figure 9. Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log and highlights the start activities (S). We plot the number of requested nodes when querying all nodes and the average number of queried nodes over all events using MFP Requesting. The diagram shows that activities with the highest average queries occur the least, excluding start activities.
[…]
[Bar diagram based on Sepsis Cases]Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log and highlights the start activities (S). The diagram shows that activities with the highest average queries occur the least, excluding start activities.
5.4.1. Comparison Across Datasets
@@
Fig. 4 and Table 1 illustrate, for each dataset, the average number of nodes queried per event to find the predecessor. Query All serves as the naive baseline of messaging every node in the network. With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event. We achieve additional message reduction when the process has a designated start event, making predecessor requests for the respective node redundant. Overall, the reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
@@
Fig. 7 shows Cumulative Distribution Functions (CDF) for all datasets. The Hospital Log has the steepest increase, with over 90% of events needing queries to no more than 5% of the nodes. Conversely, the Traffic Fine dataset has very few instances where the predecessor can be identified after querying just 5% of the nodes. This deviation stems from the limited number of 11 activities in the Traffic Fine dataset. Hence, within the 5% threshold, only self-loops are considered.
5.4.2. Stabilization over Time
@@
We now delve deeper into the specifics of the Hospital Log, as it is by far the largest one in terms of activities. Fig. 8 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events. There are still fluctuations, for example, at around 5,000 events. This event log is a good example of stabilization due to its large activity set. Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
5.4.3. Activity-wise Analysis
@@
Next, we evaluate the number of queries per activity in the Sepsis Cases log, see Fig. 9. In contrast to other event logs, the Sepsis Cases do not have one designated start event but six different ones. However, about 95% of cases start with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others. Although ”Release E” shows an increased query count, it is not a start activity. The increased query count results from the limited occurrence (only six times), indicating non converged predecessor stabilization. The situation is similar for the four activities to its left. They occur infrequently, which also explains the higher standard error.
5.4.4. Batching
@@
Fig. 7 illustrates the reduction in the average number of queried nodes per event as the batch size increases. For example, at a batch size of 40, all datasets achieve an average number of nodes queried per event of less than 2.5% of the nodes in the network. For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log, those are on average 15.6 nodes.
5.4.5. Communication Cost and Privacy
Table 2 compares the centralized approach with the naive EdgeMiner algorithm and EdgeMiner using MFP Requesting and batching in terms of communication overhead and privacy. The centralized approach requires fewer messages per event, but at the cost of reduced privacy. For the naive EdgeMiner, the number of messages per event is equal to $12\cdot(n-1)+$ 1 with $n$ the number of activities/nodes. All nodes except the node itself are contacted, all other nodes reply, and the chosen node gets a confirmation message. For EdgeMiner with MFP Requesting and batching we compute $12\cdot k+$ 1 with $k$ the average number of queried nodes per event (see Fig. 4). In the centralized approach, event data is sent to a central entity right after the event, which requires only one message per event. However, this approach inherently relies on a central storage system where all events are collected, presenting privacy and scalability risks. Although the number of messages per event for EdgeMiner with MFP Requesting and batching is higher than in the centralized approach, it is a considerable improvement compared to the naive EdgeMiner. We use a batch size of 10 events. As we show in Section 5.4.4, batching with larger batch sizes reduces the number of queried nodes and therefore the number of messages per event even further. The trade-off presented here is clear: while the EdgeMiner variants, particularly EdgeMiner with MFP Requesting and batching, cause higher communication overhead, they offer substantial scalability and privacy benefits by eliminating central storage and limiting non-aggregated communication to within the network.
Table 2. EdgeMiner’s efficiency and privacy across datasets against Centralized PM (CS/CE = Central Storage/Entity).
Centr. Naive EdgeMiner EdgeMiner with MFP & Batching (10)
Traffic Fine 1 21 1.76
Sepsis 1 31 2.32
Smart Factories 1 41 2.30
BPIC 2017 1 51 1.76
# Msg. per Event Hospital 1 1247 26.46
Storage None No CS No CS
Privacy Enablers Msg. Scope To CE In-Network In-Network
5.5. Footprint Matrix Sizes
@@
In the second phase of EdgeMiner, nodes transmit their data to a (central) entity, generating a fixed number of messages. These messages include the node’s local FM along with two additional Booleans indicating whether a node recorded the start or end activity of a trace. In EdgeMiner, each node only keeps track of their predecessors in its FM, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities. Even for datasets like the Hospital Log, which has an activity count of 624, each node transmits only a list consisting of at most 624 elements. Compared to existing distributed process discovery algorithms (Evermann and Assadipour, 2014; Gatta et al., 2019), which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size. Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources. For long-running processes, EdgeMiner can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
5.6. Evaluation Discussion
@@
EdgeMiner leverages data locality to enhance its efficiency. Unlike existing distributed process discovery algorithms (Evermann and Assadipour, 2014; Gatta et al., 2019), which typically collect the event data completely before processing it and creating FMs or process models, EdgeMiner adopts a proactive strategy by processing data at runtime and at its source. When the FM is requested, EdgeMiner has already started the process discovery algorithm, by collecting partial FMs at each node, and thereby speeding up mining. Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events. When looking at conformance checking, EdgeMiner enables online conformance checking by distributing the computation of FMs and making the partial FMs available at the data sources.
6. Conclusion
@@
In this paper, we introduce EdgeMiner, a novel approach tailored to enable distributed process discovery and conformance checking in IoT networks. By distributing computations to IoT and edge devices, EdgeMiner addresses scalability and privacy concerns inherent in centralized approaches. We discuss how EdgeMiner’s design principles support common process mining algorithms that construct footprint matrices or record direct successions, such as the Alpha, Heuristics and Inductive Miner (van der Aalst et al., 2004; Weijters et al., 2006; Leemans et al., 2013). Moreover, integrating Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network. The batching of predecessor queries further minimizes communication overhead.
Our empirical evaluations showcase EdgeMiner’s stark reduction of messages in real-world scenarios compared to a naive predecessor determination approach. EdgeMiner achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event. By batching predecessor queries in groups of, for example, 40 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes. We analytically and experimentally demonstrate the correctness of the algorithm and show how the intermediate results of EdgeMiner quickly converge to the FM of the full event log. We highlight how EdgeMiner provides a significant time advantage over algorithms lacking pre-processing capabilities while still producing the same FMs as the centralized algorithms. By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeMiner enables data privacy through aggregation that eliminates personal identification.
Acknowledgements.
This work received funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), grant 496119880.
Appendix A Appendix: Formal Correctness Proof
We introduce a formal correctness proof. For this, we show that the merging of partial FMs in EdgeMiner outputs the same FM as calculating the FM based on a full event log.
We begin by introducing notations before defining our claim. Let $FM^{all}$ denote the FM computed from an entire event log $L$ and $n:=|T_{L}|$, where $T_{L}$ denotes the set of all activities in $L$. Further, let $FM^{i}$ for $i\in\{0,\dots,n-1\}$ be the partial FM computed by node $i$, and $L_{i}$ the partial event log that node $i$ has knowledge of. Since $i$ only keeps track of its predecessor events, its footprint matrix $FM^{i}$ contains only one column. In the following correctness claim, we only consider the footprint matrices and omit start and end activities. The design of EdgeMiner ensures that a node considers an activity only a start activity or end activity if it corresponds to the start or end event of a full trace. Further, we assume that the footprint matrices only consist of ones and zeroes. The correctness for footprint matrices with entries in $\mathbb{N}_{0}$ follows from the fact that EdgeMiner detects all direct successions and adds a direct succession to exactly one partial footprint matrix.
Claim 1.
Proof.
It remains to show that the opposite direction holds. Let $k>_{L}\ell,~{}k,\ell\in\{0,\dots,n-1\}$ be a direct succession represented in $FM^{all}$, formally $FM^{all}_{k\ell}=1$. Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$. Merging all partial FMs only means concatenating them ordered by activity ID. Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$. Thus proving the claim. ∎